OpenStudy (anonymous):

What is the length of the hypotenuse of the triangle with the bottom leg 5 square root 2 the right leg 5 square root 2 and hypotenuse is h?

OpenStudy (tanya123):

In any right triangle, the hypotense squared = the square of one side + the square of the other side. h2 = a2 + b2 Here, h2 = 52 + 122 = 25 + 144 = 169 We have h2 = 169 The square root of 169 is 13, [ so the length of the hypotenuse is 13 inches. length of hypotenuse = 13 ]

OpenStudy (anonymous):

Use the pythagorean theorem here: The square of the hypotenuse is equal to the sum of the squares of the sides. This is often written a^2 + b^2 = c^2

OpenStudy (anonymous):

\[5\sqrt{2}^{2}+5\sqrt{2}^{2}=h ^{2}\]

OpenStudy (anonymous):

i need to know what this is quick >.<

OpenStudy (anonymous):

im sorry to confuse you but that was just a texting face sorry for the confusion. can you help me with the question i just put up though?

OpenStudy (tanya123):

lol... i didnt pay attension to the face !!! read the first post i explained to the whole thing !

OpenStudy (anonymous):

that isnt one of the choices for multiple choice question im sorry if i come off as lazy but im graduating tomorrow and if i dont finish this class by tomorrow i will not graduate and i need all the assistance i can get

OpenStudy (tanya123):

oh.. sorry about that :(

OpenStudy (tanya123):

wht class is it exactly ???

OpenStudy (anonymous):

its geometry but its an online class called apex and im so close to finishing but this is the last section and im drawing blanks on the questions

OpenStudy (anonymous):

The way I read your problem, the sides are \[5\sqrt{2}\] The square of that isn't \[5 \sqrt{2}^2\] it's \[(5\sqrt{2})^2\]

OpenStudy (anonymous):

can i send you a message of the actual problem? and possible answers?

OpenStudy (anonymous):

Let's see if I have your problem deciphered correctly: What is the square root of 2?

OpenStudy (anonymous):



OpenStudy (anonymous):

this is what the question looks like

OpenStudy (anonymous):

Pretty picture. Now, what is the square root of 2?

OpenStudy (tanya123):

the square root of 2 is 1.4142135623730950488016997242097

OpenStudy (anonymous):

Correct. Now what is 5 times the square root of 2?

OpenStudy (anonymous):

7.071067811865475

OpenStudy (tanya123):

7.0710678118654752440084436210485

OpenStudy (anonymous):

Also correct. And THAT is the length of 2 of your sides. Take that 7.071 and square it. Then double it. That effectively gives you: 7.071^2 + 7.071^2 = the square of the hypotenuse.

OpenStudy (anonymous):

so the answers 10?

OpenStudy (anonymous):

YES!!!!! EXCELLENT!!! :-)

OpenStudy (anonymous):

7.071^2 = 50 50+50=100 sqrt(100)=10

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (anonymous):

would you still be able to help me?

OpenStudy (anonymous):

No problem! I'm glad I was able to walk you through it! :-)

OpenStudy (anonymous):

With what? I have to get off here pretty soon now.

OpenStudy (anonymous):

And thank you! :-)

OpenStudy (tanya123):

@qweqwe123123123123111 u r an expert !!!

OpenStudy (anonymous):

Only at the stuff I know... With stuff I don't know, I'm just as big an idiot as anyone else... :-)

OpenStudy (anonymous):

 what is the length of the diagonal of the square?

OpenStudy (tanya123):

lol... soo ur a geometry genius ???

OpenStudy (anonymous):

Okay, this one's an easy quicky... it's the sqrt(11^2 + 11^2) It's done exactly the same way as your last problem.

OpenStudy (anonymous):

Nah. I'm a hypotenuse expert. :-)

OpenStudy (anonymous):

is it \[11\sqrt{2} ?\]

OpenStudy (anonymous):

No, it's \[\sqrt{11^{2}+11^{2}}\]

OpenStudy (anonymous):

but that equals 15.55634918610405 and wouldnt that be equivalent to \[11\sqrt{2}\]?

OpenStudy (anonymous):

Gakk!! You're getting ahead of me!! Yes, of course it is. I wasn't looking at it that way, but it is. Good job! :-)

OpenStudy (anonymous):

its all cause of you! thanks a ton

OpenStudy (anonymous):

No problem at all! :-)

OpenStudy (tanya123):

Good luck for ur graduation !!! do inform !

OpenStudy (anonymous):

i may be asking a few other questions through out the night so please if you can help id VERY MUCH appreciate it tanya

OpenStudy (tanya123):

i will try ... :/